OpenStudy (anonymous):

Please help...fan and medal!! Solve the equation. |k + 9| = 4 A. {–13, –5} B. {–5, 5} C. {–5, 13} D. {all real numbers greater than or equal to –13 and less than or equal to –5}

OpenStudy (anonymous):

solve the piecewise k+9=4 and -(k+9)=4

OpenStudy (anonymous):

its B because -5+9=4

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \bf k + 9 =± 4}\) solve each

OpenStudy (anonymous):

@thing_one Where is thing 2?

OpenStudy (anonymous):

lel

OpenStudy (anonymous):

@SolomonZelman and @thing_one those aren't the correct methods

OpenStudy (anonymous):

@AnswerMyQuestions im not sure he might be lost

OpenStudy (anonymous):

or actually @SolomonZelman you're correct but @thing_one isn't

OpenStudy (anonymous):

I think he's with the cat in the hat.

OpenStudy (neer2890):

If k is an integer then it should be [-13,-5] if k is not an integer then it should be d option

OpenStudy (anonymous):

yeah i was fired from that so i decided to answer questions instead of being a pain